OpenStudy (anonymous):

The function below shows the number of students of a school who enrolled for yoga classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 15(1.23)x The average rate of change in the number of students who enrolled for yoga classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = 15(1.23)^x }\) like this right?

OpenStudy (anonymous):

Yup

OpenStudy (solomonzelman):

3rd year. \(\normalsize\color{blue}{ f(x) = 15(1.23)^3≈27.91 }\) 5th year. \(\normalsize\color{blue}{ f(x) = 15(1.23)^5≈42.23 }\) \(\normalsize\color{blue}{ (42.23-27.91) \div 2. }\) \(\normalsize\color{blue}{ 14.32 \div 2= }\) \(\normalsize\color{blue}{ \underline{7.16 } }\) (answer)

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Anytime, enjoy !